The manager of a gas station decided to identify his customers' demand curve for gasoline of in order to improve the profitability of his gas station. For this reason he collected data during a three months interval of time and on this basis created the following table indicating the relationship between the price of a gallon of gasoline and the average number of gallons sold per day:

Price    Number of gallons

$2.09 2143

$2.19 2081

$2.29 2097

$2.39 2006

$2.49 2011

$2.59 1987

$2.69 1937

$2.78 1901

$2.89    1915

$2.99    1924

$3.09    1870   

$3.19    1822

$3.29 1767

$3.39    1725

$3.49. 1693

$3.59    1681

Taking into consideration these data,

1. present them graphically by using EXCEL Graphics

2. identify the linear demand function by using EXCEL Statistical Functions

3. determine the hypothetical price for which the demand might be zero

4. determine the size of the demand in the hypothetical case that the gasoline would be free (price = zero dollars!!!)

5. analyze the relevance and validity of a linear demand curve taking into consideration the #3 and #4 results.
Answer #1
Price Number of gallons
$2.09 2143
$2.19 2081
$2.29 2097
$2.39 2006
$2.49 2011
$2.59 1987
$2.69 1937
$2.78 1901
$2.89 1915
$2.99 1924
$3.09 1870
$3.19 1822
$3.29 1767
$3.39 1725
$3.49 1693
$3.59 1681

graphically by using EXCEL Graphics

Demand Curve 2500 y=-29.632x + 2161.9 R = 0.968 2000 Number of Gallons $3.19 $3.29 $3.39 $3.49 $3.59 Price

the linear demand function is written on the above graphics

the hypothetical price for which the demand might be zero and the size of the demand in the hypothetical case that the gasoline would be free (price = zero dollars!!!)

x=Price y ==29-632 y=quantity. +2161-9 il x=0 & Price=0 Y = -29.632 (x) + 216109 Y = 216109 if y=0 s quantity=0 then o=-29-63

the relevance and validity of a linear demand curve taking into consideration the #3 and #4 results.

Demand Curve 2500 2000 Quantity 72.95828830 Price -Demand Curve